OpenStudy (anonymous):

What is the radius of the circle (x-2)^2+(y+3)^2=6

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf (x-h)^2+(y-k)^2=r^2 }\)

OpenStudy (solomonzelman):

(h,K) is the center. r is the radius. Compare your equation to the blue.

OpenStudy (anonymous):

So is the radius 6 or 36?

OpenStudy (solomonzelman):

but in your case, from the comparison, you can see that \(\large\color{blue}{ \bf 6=r^2 }\)

OpenStudy (solomonzelman):

So r = ?

OpenStudy (anonymous):

6

OpenStudy (solomonzelman):

close r^2=6 r=±√6 but in your case just √6 (and not -√6 ) because radius is distance and can't be negative.

OpenStudy (anonymous):

Ohhhh okay thank you! Since you are here can you really quick help me out with another. Find the center and the radius of the circle: x^2+y^2-12x+6y=19

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

x^2+y^2-12x+6y=19 x^2-12x+y^2+6y=19 x^2-12x+36+y^2+6y=19+36 x^2-12x+36+y^2+6y+9=19+36+9 (x-6)^2+(y+3)^2=64 (x-6)^2+(y+3)^2=8^2

OpenStudy (solomonzelman):

compare to (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

Is the answer center (-6,3) and radius 64?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

the other way (6,-3) and the radius is 8

OpenStudy (anonymous):

Thank you so much!